OpenStudy (anonymous):

2, -8, 32, -128, 512

OpenStudy (freckles):

hi!

OpenStudy (freckles):

think about this for a second: 2^1=2 2^2=4 2^3=8 2^4=16 2^5=32 2^6=64 2^7=128 look at when the power is odd what do you notice about your sequence of numbers ?

OpenStudy (freckles):

then also consider that your sequence alternates